The sum of two numbers is 5/4. One number is 1/2 of the other. Find the
product of these two numbers.
A. 25/72; B. 25/63; C. 5/8; D. 5/6; E. none

Answer by Theo(2973)

x + y = 5/4

x = .5*y

this makes:

y = 2*x

substituting in original equation gets:

x + 2*x = 5/4

combine like terms to get:

3*x = 5/4

divide both sides by 3 to get:

x = 5/12

y = 2x so:

y = 5/6

substitute in original equation of x + y = 5/12 to get:

5/12 + 5/6 = 5/12 + 10/12 = 15/12 = 5/4

since x = 5/12 and y = 5/6, the product of the two numbers is:

5/12 * 5/6 = 25/72

that would be selection A.